OpenStudy (anonymous):

what equation in slope intercept form represents the line that passes through the points 4,5 and 6,2

jigglypuff314 (jigglypuff314):

slope intercept form: y = mx + b where m=slope and b=y-intercept so given points (4, 5) and (6, 2) first find the slope between then given \[slope = \frac{ rise }{ run } = \frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]when given points (x1, y1) and (x2, y2)

OpenStudy (anonymous):

i dont follow

jigglypuff314 (jigglypuff314):

first find the slope between the two points :)